OpenStudy (aurange):

If y = 1 + i, then y3 − 3y2 + (____) − 1 = -i . pls help

OpenStudy (amorfide):

replace the space with a variable \[y^{3}-3y^{2}+x-1=-i\] y=1+i substitute it in \[(1+i)^{3}-3(1+i)^{2} + x - 1 = -i\] and solve for x